OpenStudy (anonymous):

Find the sine, cosine, and tangent of 45 degrees. Sin 45 degrees = negative square root of 2 divided by 2, cos 45 degrees = negative square root of 2 divided by 2, tan 45 degrees = negative square root of 2 Sin 45 degrees = square root of 2 divided by 2, cos 45 degrees = square root of 2 divided by 2, tan 45 degrees = square root of 2 Sin 45 degrees = square root of 2 divided by 2, cos 45 degrees = square root of 2 divided by 2, tan 45 degrees = 1 Sin 45 degrees = square root of 2 divided by 2, cos 45 degrees = square root of 2 divided by 2, tan 45 degrees = −1

OpenStudy (johnweldon1993):

Remember your special triangles, it can make a problem like this very quick A 45-45-90 triangle 

OpenStudy (johnweldon1993):

\[\large sin(\theta) = \frac{opposite}{hypotenuse}\] \[\large cos(\theta) = \frac{adjacent}{hypotenuse}\] \[\large tan(\theta) = \frac{opposite}{adjacent}\]

OpenStudy (anonymous):

oh now i remember!!! thank you so mcuh!

OpenStudy (johnweldon1993):

No problem :)